Cost Accounting =

Find a numerical example of cost-volume-profit (CVP) analysis, and analyze how CVP analysis is used for decision making?
Answer #1

Find a numerical example of cost-volume-profit (CVP) analysis, and analyze how CVP analysis is used for decision making?

CVP analysis

It is a technique that helps to identify the effect of sales volume and product cost on the profit of a business. It analyses the sales price, fixed costs, variable cost and the number of goods sold and measure how it affects the profit of a firm. It facilitates management to examine the relationship between cost and revenue of a firm.

COST VOLUME PROFIT ANALYSIS EXAMPLES 1

ABC ltd wishes to make an annual profit of $100000 from the sale of appliances. Details of manufacturing and annual capacity are as follows:

Description: Cost Volume Profit Analysis Examples1

Based on the above information let’s plug the numbers in CVP equation:

  • 10000*p= (10000*30) +$30000+$100000
  • 10000p = ($300000+$30000+$100000)
  • 10000p=$430000
  • Price per unit= ($430000/10000) = $43

Thus price per unit comes out to $43 which implies that will have to price its product $43 and need to sell 10000 units to achieve its targeted profit of $100000. Further, we can see that the fixed cost remain constant ($30000) irrespective of the level of sales.

COST VOLUME PROFIT ANALYSIS EXAMPLES 2

ABC Limited has entered into the business of making Electrical fans. The management of the company is interested in knowing the breakeven point at which there will be no profit/loss. Below are the details pertaining to the cost incurred:

Description: Cost Volume Profit Analysis Examples2

No. of units sold by ABC limited: ($300000/$300) = 1000 units

Variable cost per unit= ($240000/1000)=$240

  • Contribution per unit= Selling price per unit-Variable Cost per unit
  • = ($300-$240)
  • = $60 per unit

Break Even Point= (Fixed Cost/Variable Cost per unit)

  • = ($60000/$60)
  • =10000 units

Thus ABC limited need to sell 10000 units of electric fans to break even at the current cost structure.

Decision Analysis

  1. CVP analysis facilitates the decision maker to find out the break even level of sales where no profit and no loss can be achieved by a firm. In the Example, ABC limited need to sell 10000 units of electric fans to break even at the current cost structure.
  2. CVP analysis helps management to understand the different cost such as variable cost and fixed cost at different levels of production/sales volume. CVP analysis helps decision makers in forecasting cost and profit on account of change in volume.
  1. It makes us to understand the Effects of changes in fixed and variable cost help management decide the optimum level of production.